A rs 5000 par value bond bearing a coupon rate of 10% will mature after 10 years. What is the price to buy the bond at today , if discounted rate is 12%


1
Expert's answer
2021-05-19T11:00:47-0400

Bond Price = ∑(C/ (1+YTM)n )+ P / (1+i)n

Where

  • n = Period which takes values from 0 to the nth period till the cash flows ending period
  • C= Coupon payment in the nth period
  • YTM or i= interest rate or required yield
  • P = Par Value of the bond

"n=10years"

Cn=10% of 5000 = 500

YTM or i = 12%= 0.12

P= 5000


"1+Interest = 1+\frac{12}{100}= 1.12"



"Bond Price =\frac{500}{(1.12)}
+\frac{500}{(1.12)^{2}}+\frac{500}{(1.12)^{3}}+ \frac{500}{(1.12)^{4}}+ + \frac{500}{(1.12)^{5}}
+ \frac{500}{(1.12)^{6}} + \frac{500}{(1.12)^{7}}
+ \frac{500}{(1.12)^{8}}+ + \frac{500}{(1.12)^{9}}+
\frac{500}{(1.12)^{10}}+\frac{5000}{(1.12)^{10}}"


"BondPrice = 446.43+398.60+ 345.90+317.76+283.71+253.32+226.17+201.94+170.31+160.99 + 1609.87=4415"